alessia is paid 1.4 times her normal hourly rate for each hour she works over 30 hours in a week. last week she worked 35 hours and earned $436.60. what is her normal hourly rate?
Answer by checkley77(12569)

30x+5*1.4x=436.60
30x+7x=436.60
37x=436.60
x=436.60/37
x=$11.80 is the normal hourly rate.
1.4*11.80=$15.52 is the overtime horly rate.
Proof:
30*11.80+5*15.52=436.60
354+82.60436.60436.60=436.60